# Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs

### v3

@@ 1. Introduction
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where bsbs is the batch size, slsl sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (wijmans2024cut). However, in SRS, $|V|$ can reach millions of items (zivic2024scaling), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (touvron2023llama), QWEN (bai2023qwen), and Gemma-7B (team2403gemma), respectively.
@@
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (kang2018self) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(sun2019bert4rec; klenitskiy2023turning). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (klenitskiy2023turning; mezentsev2024scalable). Instead of scoring all items during training, a subset of negative samples (nsns) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing nsns in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (di2024theoretical; klenitskiy2023turning; guo2024scaling). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (mezentsev2024scalable; klenitskiy2023turning) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(prakash2024evaluating). This creates an open question of whether practitioners should maximize nsns given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
@@
    Comprehensive Memory-Accuracy Trade-off Analysis: To address Q1, we conducted a systematic investigation into the relationships among accuracy of SASRec, batch size (bsbs), sequence length (slsl), the number of negative sample size (nsns) across six datasets with varying item catalog sizes ($|V|$). Our findings indicate that simply maximizing nsns in $\mathrm{CE}^{-}$ or employing CE to encompass the entire item catalog $|V|$ leads to suboptimal performance in memory-constrained environments. Instead, the choice between $\mathrm{CE}^{-}$ and CE requires a data set-specific calibration. We provide analysis for configuring these parameters to balance accuracy and memory efficiency.
    Furthermore, we make publicly available a dataset comprising 997 data points, each representing SASRec trained across the considered datasets. For every configuration, we log bsbs, slsl, nsns, memory consumption, and various metrics along the training steps. This dataset enables reproducible analysis of memory-accuracy trade-offs.
@@
    Memory-Efficient Metric Improvement: By reallocating the memory savings achieved through the use of CCE and $\mathrm{CCE}^{-}$ to increase batch size bsbs, sequence length slsl, and the number of negative samples nsns, we attain up to a 30% improvement in accuracy while maintaining the same computational budget.
@@ 3.1. Cut-Cross-Entropy for RecSys
In this study, we incorporated CCE into sequential recommendation transformer models to evaluate its effectiveness for recommendation tasks. For this purpose, the three-dimensional output tensor of embeddings produced by the final transformer block is flattened into a matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ represents the hidden size and NN corresponds to the total number of items, calculated as the product of the batch size bsbs and the sequence length slsl. The CE loss for the set of items of size |Vvert|V\ vert is written as follows:
@@ 3.2. Sparse Gradients in Training with CE loss
$3.$ Cut-Cross-Entropy with Negative Sampling
@@
$\mathrm{CCE}^{-}$ Loss Formulation. The loss function operates on an embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, which is obtained by flattening the 3D output tensor along its first two dimensions. During training, at each time step, the positive items corresponding to users are augmented with a dynamically sampled set of nsns negative items. This creates a compact item subset $V_{s}\subset V$ where: $|V_{s}|=1+ns$ (1 positive + nsns negatives) and $|V_{s}|\ll|V|$ (sampled subset size $\ll$ size of the full catalog of items ). The loss calculation is thereby confined to interactions within this compact subset, enabling efficient training, as discussed below. The $\mathrm{CCE}^{-}$loss is written as
@@ 5. Results
In the following, we present the results of our computational experiments and analyze them to address the questions Q1 and Q2 posed in Section 1 and evaluate the performance of CCE and $\mathrm{CCE}^{-}$. Specifically, (1) Section 5.1 contains the results of experiments carried out to study the effect of training hyperparameters on SASRec performance, where, among other insights, we demonstrate that sometimes increasing nsns does not lead to an improvement in the accuracy of the model; therefore, sometimes $\mathrm{CE}^{-}$ should be preferred over CE. (2) Further, we compare the training efficiency of the SASRec model using CCE and $\mathrm{CCE}^{-}$ over CE, $\mathrm{CE}^{-}$ in Sections 5.2. The results of this section further highlight the effectiveness of negative sampling in SRS. (3) Since uniform negative sampling is not the only possible sampling strategy in Section 5.4.1, we study other negative sampling strategies. (4) Finally, we extend our empirical analysis to the Bert4Rec model in Section 5.6.
@@ 5.1. Optimal Accuracy-Memory Trade-off
Figure 6. This figure presents aggregated results from Figure 5, averaging across sequence lengths. Error bars indicate standard deviations. Colors represent batch sizes (bsbs). Performance increases with negative sample size (nsns) for some datasets, while saturating for others (Movielens-20M, Gowalla). Batch size scaling shows no universal pattern: larger batches generally improve performance but saturate for 30Music, while smaller batches (128 vs 512) yield better results for Movielens-20M
@@
Table 3. Comparison of CE and CCE. Hyperparameters such as $bs,sl$, and nsns were chosen based on our initial pool of scaling experiments for CE and $\mathrm{CE}^{-}$, respectively. CCE and $\mathrm{CCE}^{-}$ were trained using the same settings. We report NDCG@10, allocated memory (GB) and time per epoch (seconds). Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to CE and $\mathrm{CE}^{-}$, green denotes CCE and $\mathrm{CCE}^{-}$ superiority.
@@
Table 4. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except nsns for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 5. Comparison Coverage@10 and Surprisal@10 time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except nsns for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@ 5.3. Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
The results discussed in the previous section indicate that CCE and $\mathrm{CCE}^{-}$ significantly reduce the memory consumption required for training SASRec. This reduction provides an opportunity to enhance model performance by increasing batch size (bsbs), sequence length (slsl), and number of samples (nsns) in the case of $\mathrm{CCE}^{-}$, implementing insights from the Section 5.1. In these experiments, we set identical training hyperparameters for both CCE and $\mathrm{CCE}^{-}$. Tables 4, 5 present the values of the NDSG@10, Coverage@10, and Surprisal@10 metrics for all datasets, along with the allocated memory and time per epoch required for training.
@@ 5.4.1. Comparison Sampling Strategies
$5.$ Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE

### v4

@@ 1. Introduction
Peak memory is one of the main computational constraints when cross-entropy loss (CE) is applied to large catalogs. The memory footprint at the final layer for logits can be estimated as $bs\cdot sl\cdot|V|$ [^1] To illustrate, consider a batch of 256 sequences, each of length 100, with a catalog containing $10^{6}$ items. The resulting logit tensor would require the storage of $256\cdot 100\cdot 10^{6}$ elements, equating to approximately 100 GB of memory., where $bs$ is the batch size, $sl$ sequence length and $|V|$ is the catalog or vocabulary size, typically the most significant contributor. This final layer uses up to 90% of the combined memory footprint, gradients, and activations, when training modern Large Language Models (LLMs) (wijmans2024cut). However, in SRS, $|V|$ can reach millions of items (zivic2024scaling), surpassing even LLMs, which have vocabulary sizes orders of magnitude smaller: 32K, 152K, and 256K for Llama2-7B (touvron2023llama), QWEN (bai2023qwen), and Gemma-7B (team2403gemma), respectively.
@@
This memory footprint has been a key driver for early SRS adopting the binary cross-entropy loss (BCE) (kang2018self) and the cross-entropy loss with negative sampling ($\mathrm{CE}^{-}$(sun2019bert4rec; klenitskiy2023turning). $\mathrm{CE}^{-}$ directly addresses memory constraints caused by large item catalogs (klenitskiy2023turning; mezentsev2024scalable). Instead of scoring all items during training, a subset of negative samples ($ns$) is used, where $ns\ll|V|$, to approximate the full CE loss. This reduces computational complexity and memory requirements but introduces a hyperparameter trade-off between batch size, sequence length, and number of negative samples (ns) in memory-constrained scenarios. Prior work indicates that maximizing $ns$ in $\mathrm{CE}^{-}$ or using CE can improve performance of the models (di2024theoretical; klenitskiy2023turning; guo2024scaling). However, empirical results demonstrate that $\mathrm{CE}^{-}$ can sometimes outperform full CE for some datasets (mezentsev2024scalable; klenitskiy2023turning) and the performance of models trained with $\mathrm{CE}^{-}$ shows a non-linear dependence on nsns(prakash2024evaluating). This creates an open question of whether practitioners should maximize $ns$ given a fixed memory budget or prioritize other memory-related parameters. We will refer to this question as the first question Q1.
@@
    Comprehensive Memory-Accuracy Trade-off Analysis: To address Q1, we conducted a systematic investigation into the relationships among accuracy of SASRec, batch size ($bs$), sequence length ($sl$), the number of negative sample size ($ns$) across six datasets with varying item catalog sizes ($|V|$). Our findings indicate that simply maximizing $ns$ in $\mathrm{CE}^{-}$ or employing CE to encompass the entire item catalog $|V|$ leads to suboptimal performance in memory-constrained environments. Instead, the choice between $\mathrm{CE}^{-}$ and CE requires a data set-specific calibration. We provide analysis for configuring these parameters to balance accuracy and memory efficiency.
    Furthermore, we make publicly available a dataset comprising 997 data points, each representing SASRec trained across the considered datasets. For every configuration, we log $bs$, slsl, nsns, memory consumption, and various metrics along the training steps. This dataset enables reproducible analysis of memory-accuracy trade-offs.
@@
    Memory-Efficient Metric Improvement: By reallocating the memory savings achieved through the use of CCE and $\mathrm{CCE}^{-}$ to increase batch size $bs$, sequence length $sl$, and the number of negative samples $ns$, we attain up to a 30% improvement in accuracy while maintaining the same computational budget.
@@ 3.1. Cut-Cross-Entropy for RecSys
In this study, we incorporated CCE into sequential recommendation transformer models to evaluate its effectiveness for recommendation tasks. For this purpose, the three-dimensional output tensor of embeddings produced by the final transformer block is flattened into a matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ represents the hidden size and $N$ corresponds to the total number of items, calculated as the product of the batch size $bs$ and the sequence length $sl$. The CE loss for the set of items of size $|V|$ is written as follows:
@@ 3.2. Sparse Gradients in Training with CE loss
3.3. Cut-Cross-Entropy with Negative Sampling
@@
$\mathrm{CCE}^{-}$ Loss Formulation. The loss function operates on an embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, which is obtained by flattening the 3D output tensor along its first two dimensions. During training, at each time step, the positive items corresponding to users are augmented with a dynamically sampled set of $ns$ negative items. This creates a compact item subset $V_{s}\subset V$ where: $|V_{s}|=1+ns$ (1 positive + $ns$ negatives) and $|V_{s}|\ll|V|$ (sampled subset size $\ll$ size of the full catalog of items ). The loss calculation is thereby confined to interactions within this compact subset, enabling efficient training, as discussed below. The $\mathrm{CCE}^{-}$loss is written as
@@ 5. Results
In the following, we present the results of our computational experiments and analyze them to address the questions Q1 and Q2 posed in Section 1 and evaluate the performance of CCE and $\mathrm{CCE}^{-}$. Specifically, (1) Section 5.1 contains the results of experiments carried out to study the effect of training hyperparameters on SASRec performance, where, among other insights, we demonstrate that sometimes increasing $ns$ does not lead to an improvement in the accuracy of the model; therefore, sometimes $\mathrm{CE}^{-}$ should be preferred over CE. (2) Further, we compare the training efficiency of the SASRec model using CCE and $\mathrm{CCE}^{-}$ over CE, $\mathrm{CE}^{-}$ in Sections 5.2. The results of this section further highlight the effectiveness of negative sampling in SRS. (3) Since uniform negative sampling is not the only possible sampling strategy in Section 5.4.1, we study other negative sampling strategies. (4) Finally, we extend our empirical analysis to the Bert4Rec model in Section 5.6.
@@ 5.1. Optimal Accuracy-Memory Trade-off
Figure 6. This figure presents aggregated results from Figure 5, averaging across sequence lengths. Error bars indicate standard deviations. Colors represent batch sizes ($bs$). Performance increases with negative sample size ($ns$) for some datasets, while saturating for others (Movielens-20M, Gowalla). Batch size scaling shows no universal pattern: larger batches generally improve performance but saturate for 30Music, while smaller batches (128 vs 512) yield better results for Movielens-20M
@@
Table 3. Comparison of CE and CCE. Hyperparameters such as $bs,sl$, and $ns$ were chosen based on our initial pool of scaling experiments for CE and $\mathrm{CE}^{-}$, respectively. CCE and $\mathrm{CCE}^{-}$ were trained using the same settings. We report NDCG@10, allocated memory (GB) and time per epoch (seconds). Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to CE and $\mathrm{CE}^{-}$, green denotes CCE and $\mathrm{CCE}^{-}$ superiority.
@@
Table 4. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
Table 5. Comparison Coverage@10 and Surprisal@10 time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@ 5.3. Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
The results discussed in the previous section indicate that CCE and $\mathrm{CCE}^{-}$ significantly reduce the memory consumption required for training SASRec. This reduction provides an opportunity to enhance model performance by increasing batch size ($bs$), sequence length ($sl$), and number of samples ($ns$) in the case of $\mathrm{CCE}^{-}$, implementing insights from the Section 5.1. In these experiments, we set identical training hyperparameters for both CCE and $\mathrm{CCE}^{-}$. Tables 4, 5 present the values of the NDSG@10, Coverage@10, and Surprisal@10 metrics for all datasets, along with the allocated memory and time per epoch required for training.
@@ 5.4.1. Comparison Sampling Strategies
5.5. Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE